you travel 350 miles in the same amount of time it takes a friend to travel 400 miles. your friends average speed is 4 miles per hour greater than your average speed. find your friends average speed
Answer by mananth(16949)

My speed x mph
Friend's speed 4 faster x + 4 mph

My speed 350 miles
Friend's speed 400 miles
My time = Friend's time
t=d/r
350 / x = 400 /( x + 4 )
cross multiply
350 ( x + 4 ) = 400 x
350 x + 1400 = 400 x
350 x - 400 x = -1400
-50 x = -1400
/ -50
x = 28 mph

150 mph My speed
154 mph Friend's speed